Questions 1. Determine the empirical formula of the oxide of magnesium. (Think carefully about the precision of each mass that you measured before you do this caleulation. How many significant figures are justified fo each mass?) all Anc $8 FESEIOS IN.S itoesM bi boe ofd )bi ban sld wus to a.cm.oer e.ce
Results 33 306 Mass of empty crucible and lid (g) 33.5329 33.b4I 33.Y12 Average mass of empty crucible and lid (g) 33.513 Mass of crucible, lid, and Mg (g) 33.412 Average mass of empty crucible and lid (g) otstpo isoario beouied 0.2 olhsMass of Mg (g) noalom it ure 33.601 Mass of crucible, lid, and oxide (g) 33.595 lnelom d lew aw obi 33.598 Average mass (g) 33.412 Average mass of empty crucible and lid (g) Mass of oxide (g) 25
Answer #1

Ans:

Average mass of empty crucible + lid = (33.306 +33.290+33.641)/3 = 33.412 g

Mass of crucible + lid + magnesium (Mg) = 33.513 g

Mass of magnesium taken = (33.513 – 33.412) = 0.101 g

Average mass of crucible + lid + oxide = (33.601 + 33.595)/2 = 33.598 g

Mass of magnesium oxide = (33.598 – 33.412) = 0.186 g

Mass of oxygen in magnesium oxide = (0.186 – 0.101) = 0.085 g

Mass (g)

Atomic mass (g/mol)

Number of moles

Ratio

Mg

0.101

24.305

0.101 /24.305 = 0.0042

0.0042/0.0042 = 1

O

0.085

15.999

0.085/15.999 = 0.0053

0.0053/0.0042 = 1.3 ~1

That means for every single magnesium there is one ‘O’

So empirical formula will be Mg1O1 = MgO